A ball is dropped from a height of 64 feet. Its height above the earth in feet is given by h(t ) = -16t2 + 64, where t is the number of seconds after it is dropped. Find h(1). How long does it take for the ball to fall to the earth?
Answer by Alan3354(69443)

A ball is dropped from a height of 64 feet. Its height above the earth in feet is given by h(t ) = -16t2 + 64, where t is the number of seconds after it is dropped. Find h(1). How long does it take for the ball to fall to the earth?
-----------------
h(1) = -16 + 64 = 48 feet
--------------
Set h = 0 to find when it impacts
0 = -16t^2 + 64
t^2 = 4
t = 2 seconds (ignore the -2)